OpenStudy (anonymous):

When x is divided by 5, the remainder is 3. What is the remainder when 4x is divided by 4?

OpenStudy (anonymous):

try with a number and see

OpenStudy (anonymous):

or note that any number multiplied by 4 is a multiple of 4!

OpenStudy (anonymous):

makes no difference what the remainder is when divided by any other number, \(4x\) is a multiple of \(4\) no matter what

OpenStudy (tkhunny):

x = 5*m + 3 where m is a Whole Number and greater than 3, but we do not actually need any of this information. 4 is always divisible by 4. Thus, 0.

OpenStudy (solomonzelman):

Lets see. When x/5= remainder =3 so when 4x/5= it is 4 times 3 divide 12 by 5, you get 2.

OpenStudy (anonymous):

you are given x = 5k + 3, for some \[k \in Z\] which is also written as x = 3(mod 5) so 4x = 12(mod 5) => -x = 2(mod 5) => x = -2(mod5) => x=3 so the remainder for 4x/4 is 4x = 4k + r, for some \[k \in Z\], r = remainder 4x = r(mod 4) r = 4x(mod 4) => r = 0x(mod4) so the remainder is r = 0x, \[\forall x \in Z \]